OpenStudy (dmezzullo):

Choose the correct description of the graph of the compound inequality 3x < -9 or x + 2=> 10 A number line with a closed circle on -3, shading to the left, and an open circle on 8, shading to the right. A number line with an open circle on -3, shading to the left, and a closed circle on 8, shading to the right. A number line with an open circle on -3, a closed circle on 8, and shading in between. A number line with a closed circle on -3, an open circle on 8, and shading in between. I think it would be C.

OpenStudy (anonymous):

\[\large 3x < -9~~ or ~~ x + 2 \ge 10\]_______________________ \[3x < -9\]Divide by 3. \[x < -9 \]Open circle, shading to the left. __ \[x + 2 \ge 10\]Subtract 2. \[x \ge 8\]Closed circle, Shading to the right. --

OpenStudy (dmezzullo):

Hey @jazy did u get my message earlier?

OpenStudy (anonymous):

I actually just saw it lol